# Meta-LoRA: Meta-Learning LoRA Components for Domain-Aware ID Personalization

### v2

@@ 4.2 Experimental setup
Implementation details. We trained the Stage-1 Meta-LoRA model for 50,000 iterations, on a single A6000 GPU. For Stage-2 training, we use 375 iterations. These values are decided empirically. Additional implementation details can be found in the Supplementary Material. The code, model weights, and the evaluation dataset will be shared publicly.
Baseline models. We compare against three state-of-the-art subject-driven image generation models: InstantID [30], PhotoMaker [17], and PuLID [9]. For fairness, we align evaluation conditions where possible. Additionally, we provide comparisons against a standard rank-1 LoRA baseline, both with and without augmented training data, trained for the same number of iterations ($375$) as Meta-LoRA, as well as longer ($675$). A more detailed explanation of the baselines, including the rationale behind various details such as diffusion backbones, is given in the Supplementary Material.
@@ 4.3 Evaluation metrics and R-FaceSim
Following recent studies [25, 30, 33, 9, 27, 24], we use four key metrics to evaluate performance: CLIP-T, CLIP-I, DINO, and the newly proposed Robust Face Similarity (R-FaceSim). The details of the first three metrics are given in the Supplementary Material. The commonly used FaceSim metric aims to evaluate how well the generated images maintain the subject’s identity based on the cosine similarity between the reference and generated images. The facial embeddings are obtained using the backbone of a face recognition model. However, we observe two major limitations with it. First, facial recognition methods primarily learn to distinguish individuals, therefore, their embeddings rather tend to lack some of the fine-grained identity details, also observed by [4, 25, 14] in various contexts. Second, commonly the same reference images are used for both personalization and face similarity evaluation. However, in feed-forward personalization models, e.g. InstantID [30] and PuLID [9], we observe a bias towards generating images with the same pose and/or gaze as the reference images. This phenomenon indicates an undesirable limitation of those models. It also tends to artificially inflate the face-similarity scores, because the face recognition embeddings are sensitive to having the same pose or gaze.
To address the former issue, we provide a detailed discussion of the qualitative results. To address the latter one, we propose a novel face-similarity metric that we call Robust Face Similarity (R-FaceSim). R-FaceSim is computed by excluding the reference image (used for fine-tuning or output-conditioning); we instead compare each generated image to other real images of the same person (with different poses) and average the cosine similarity. This yields a face-similarity score not inflated by the exact reference image. We encourage future research to adopt this approach to improve the inflated scores from pose copying as acknowledged in [7].
@@ 4.4 Quantitative Results
Table 1 provides a quantitative comparison of Meta-LoRA against the previously mentioned baselines on the Meta-PHD test set. Our evaluation focuses on three key aspects: CLIP-T, CLIP-I, and DINO scores, which assess prompt adherence and visual alignment; R-FaceSim, which measures identity preservation; and the size of the training datasets.
Due to the nature of CLIP-I and DINO, models that generate images closely resembling their base model tend to achieve higher scores in these metrics. However, strong generalization capabilities in base models can result in lower CLIP-T scores when fine-tuned for identity personalization. Meanwhile, R-FaceSim increases when the reference identity is overfitted, but excessive training can degrade generalization. Existing identity personalization models struggle to balance these competing objectives. InstantID [30] and PhotoMaker [17], for instance, show a bias toward either face similarity or text/image alignment. PhotoMaker exhibits higher CLIP-T, CLIP-I, and DINO scores at the cost of identity preservation, whereas InstantID prioritizes facial similarity at the expense of generalization. PuLID [9] and our proposed Meta-LoRA, on the other hand, achieve a better balance between identity preservation and generalization.
Compared to InstantID, our model outperforms across all evaluation metrics, achieving a more balanced trade-off between prompt adherence and identity retention. Against PhotoMaker, Meta-LoRA delivers comparable results in CLIP-I, CLIP-T, and DINO metrics while exhibiting superior facial similarity. Additionally, when compared to PuLID with the FLUX.1-dev [15] base, Meta-LoRA maintains better face similarity (i.e., a higher R-FaceSim score) while exhibiting superior prompt-following ability and a stronger capacity to inject identity-specific details without distorting the overall scene (i.e., higher CLIP-T, CLIP-I, and DINO scores). Given that both models share an implementation with the same base module, we find it most logical to compare Meta-LoRA with PuLID using the FLUX.1-dev [15] variation. However, similar observations hold for other PuLID variations as well.
@@
Another key distinction lies in the amount of pretraining data and the additional training losses employed. Existing models rely on significantly larger datasets (i.e., our dataset is 0.035% of InstantID, 1,4% of PuLID, and 18.75% of PhotoMaker). Additionally, PuLID incorporate a face-centric ID loss during training and InstantID integrates facial features to its network. Despite this, Meta-LoRA achieves competitive or superior results without the use of such explicit facial losses or features, opting instead for a simpler and more efficient training strategy.
We also evaluate default Rank-1 LoRA models trained with direct and augmented inputs. At 375 iterations, these models achieve high CLIP and DINO scores but exhibit low face similarity, indicating under-fitting. In contrast, Meta-LoRA converges faster, achieving a well-balanced state at 375 iterations. Extending Rank-1 LoRA training to 625 iterations stabilizes the model if the input is augmented, yet Meta-LoRA still outperforms Rank-1 LoRA (625 iterations, plus data augmentations) across all evaluation metrics, demonstrating its efficiency and effectiveness. A more detailed comparison is given in the Supplementary Material.
@@ 4.5 Qualitative Results
Figure 3: Qualitative comparisons. Meta-LoRA shows better ID preservation while keeping prompt alignment. Meta-LoRA keeps the pose natural (1st column), promotes identities effectively (2nd and 3rd columns), and follows the given style (4th column). Details are best viewed on screen, zoomed-in.
To evaluate the visual fidelity and prompt adherence of Meta-LoRA, we conduct a qualitative comparison against several baseline methods [9, 17, 10, 30] . To enhance the rigor of our qualitative assessment of visual outcomes, we focus our presentation on female samples. This decision is predicated on the fact that Stage-1 training incorporated a larger dataset of female subjects. A more exhaustive comparative analysis of male subjects, encompassing a broader spectrum of identities, textual prompts, and base model outputs, is available in the Supplementary Material.
Figure 3 presents a qualitative comparison of different models, where each column corresponds to a specific input identity, with the reference image displayed in the first row. The subsequent rows showcase outputs generated by each method under the same prompt. Our analysis highlights key differences in identity preservation, prompt adherence, and overall image quality. First of all, Meta-LoRA demonstrates the strongest ability to maintain facial characteristics while accurately integrating the details of the given prompt. Unlike other methods, it effectively balances identity retention with scene coherence, producing visually faithful results that align well with textual descriptions. The generated images exhibit both high visual fidelity and strong textual alignment, making Meta-LoRA the most robust approach across diverse scenarios.
In contrast, LoRA shows success in following textual prompts but struggles with preserving identity, leading to inconsistencies in facial features. InstantID, on the other hand, exhibits a strong replication of pose and composition of the reference image. While this allows generated images to be more physically consistent for a subset of prompts, it significantly compromises the model’s ability to adhere to prompts that specify different poses, viewpoints, or artistic styles requiring a change in facial geometry. This behavior explains the artificially inflated identity similarity scores under naive evaluation metrics, contrasted with its weaker performance under metrics that prioritize prompt alignment.
PhotoMaker demonstrates a reasonable ability to capture the general appearance of the subject. However, it frequently struggles to reproduce fine-grained facial details, particularly when the prompt necessitates significant alterations in context or style. Meanwhile, PuLID achieves comparable facial similarity to Meta-LoRA but tends to ”correct” or smooth out distinctive identity cues that conventional face similarity metrics may not fully capture [4]. In contrast, Meta-LoRA inherits and propagates these fine details, ensuring a more faithful representation of the reference identity. Additionally, Meta-LoRA exhibits superior prompt adherence, generating images with richer contextual elements, more detailed backgrounds, and an overall enhanced sense of naturalness.
Figure 4: Qualitative comparisons across different seeds. For each different seed, Meta-LoRA successfully generates distinct variations, while also preserving the prompt alignment. Details are best viewed on screen, zoomed-in.
Figure 4 further examines the impact of random seed variation on the generated outputs. Each column presents images generated using the same seed for LoRA, PuLID, and Meta-LoRA, allowing for a direct comparison of their consistency and adaptability. This comparison also reveals several key insights: while LoRA produces outputs that often share similarities in pose and background with those of Meta-LoRA, it tends to overfit to background elements unrelated to the target identity. This results in spurious correlations, suggesting that the model is learning incidental features from the reference image rather than focusing exclusively on core identity characteristics.
PuLID exhibits a relatively low degree of variation across different seeds, suggesting a potential limitation in its ability to explore the full range of possible outputs consistent with the prompt and identity. Its prompt following capabilities also fall short in some cases in terms of modifying facial features according to the prompt.
In contrast, Meta-LoRA achieves a desirable balance between seed-dependent variation and identity consistency. It generates diverse outputs while maintaining a high degree of fidelity to the input identity, ensuring both adaptability and accuracy across different scenarios.
@@
Overall, the qualitative results strongly corroborate our quantitative findings. Meta-LoRA consistently produces identity-accurate images across a wide range of prompts, preserving the subject’s distinctive facial features while generating scenes that closely align with textual descriptions. Additionally, it maintains meaningful variation across different seeds, offering both creativity and faithfulness—an improvement that is particularly valuable in applications requiring both identity preservation and prompt adherence.
@@ 5 Conclusion
In this paper, we present Meta-Low-Rank Adaptation (Meta-LoRA), a versatile three-layer LoRA architecture designed to enhance identity personalization in text-to-image generative models. Meta-LoRA employs a two-stage training strategy: the first stage learns generic LoRA Meta-Down layers applicable across multiple identities, while the second stage fine-tunes the LoRA Mid and LoRA Up layers for a specific identity using only a single reference image. By integrating domain-specific priors and optimizing feature learning across diverse identities, Meta-LoRA significantly improves identity preservation, robustness, and adaptability—all while maintaining computational efficiency.
To evaluate our approach, we introduce Meta-PHD, a novel identity personalization dataset used exclusively for testing, and assess Meta-LoRA’s performance using this dataset. Our results demonstrate that Meta-LoRA achieves state-of-the-art (SOTA) performance or performs comparably to existing SOTA methods. Moreover, qualitative evaluations highlight Meta-LoRA’s ability to generate identity-personalized images while effectively maintaining prompt adherence.
Future research directions include the incorporation of recent LoRA variants [8, 1, 16, 35, 12] into the Meta-LoRA framework and the exploration of the proposed method’s applicability to other model customization tasks.
@@ Acknowledgments
Supplementary Material
6 Overview
This document provides various additional details and discussions as an extension of the main Meta-LoRA paper.
7 Implementation Details
Meta-LoRA is designed using the ai-toolkit [^1] Link: https://github.com/ostris/ai-toolkit framework. Its training and evaluation are conducted on a single NVIDIA RTX A6000 GPU with 48 GB of VRAM. Under these configurations, Stage-1 training takes approximately 7 days (13 seconds per iteration, 50,000 iterations in total), while Stage-2 training lasts around 18 minutes (2.8 seconds per iteration, 375 iterations in total). For all training runs, we set the learning rate to $0.0004$, total number of stage-1 iterations ($q_{total}$) to 50,000, $q_{bucket}$ to $2,500$, $q_{warm-up}$ to $1,000$, and total number of stage-2 iterations ($q_{st2}$) to $375$. The batch sizes are selected as $4$ and $1$ in Stage-1 and Stage-2, respectively. Our model is trained using the default L2-loss and AdamW [19] is preferred as the optimizer.
8 Meta-training and Personalization Algorithms
@@
9 Details on Evaluation Metrics
Below, we give a summary of the commonly used evaluation metrics.
CLIP-T. CLIP-T measures how well the generated image aligns with the textual prompt. It is computed as the cosine similarity between the CLIP embeddings of the input prompt and the generated image. We use the CLIP ViT-B/32 model [22] to obtain the embeddings. A higher CLIP-T score indicates stronger prompt adherence.
CLIP-I. CLIP-I evaluates image consistency by comparing the output generated with personalization against that generated without personalization. The cosine similarity between these two CLIP embeddings (again computed using the ViT-B/32 variant) reflects the model’s ability to inject identity-specific information without distorting the overall scene.
DINO. The DINO metric leverages the CLS token embedding from the DINO-pretrained ViT-S/16 (following [24] implementation) backbone to compute the cosine similarity between the generated personalized image and a base image. This metric provides a measure of global image similarity that complements the CLIP-based evaluations, capturing subtle visual details and overall subject fidelity.
Figure 5: This figure visualizes the changes in the metric results of Meta-LoRA and default LoRA models with respect to number of iterations in the fine-tuning stage (stage-2 for MetaLoRA).
250250250250375375375375500500500500625625625625750750750750303030303131313132323232# of IterationsCosine Similarity (%)CLIP-T Results25025025025037537537537550050050050062562562562575075075075070707070757575758080808085858585# of IterationsCLIP-I Results2502502502503753753753755005005005006256256256257507507507507575757580808080# of IterationsCosine Similarity (%)DINO Results25025025025037537537537550050050050062562562562575075075075050505050606060607070707080808080# of IterationsR-FaceSim ResultsMeta-LoRALoRALoRA + augmentation
Conventional FaceSim details. Conventional FaceSim metric is computed by extracting facial embeddings from the generated images using an Inception-ResNet [29] model pre-trained on VGGFace2 [2], followed by calculating the cosine similarity between these embeddings and those of the real subject. The a discussion on the shortcomings of the conventional FaceSim metric, and the proposed improved version is given in the experiments section of the main paper.
10 LoRA and Meta-LoRA Training Curves
In this section, we analyze the performance dynamics of both LoRA and Meta-LoRA models during fine-tuning. Evaluations are conducted across the four primary metrics used in this paper—CLIP-T, CLIP-I, DINO, and R-FaceSim—at specific iteration points: $q\in\{250,375,500,625,750\}$. The results are presented in Figure 5.
As illustrated in the figure, Meta-LoRA exhibits a faster increase in R-FaceSim and a quicker decline in the other metrics compared to LoRA, indicating that it reaches a balanced state earlier, around 375 iterations. In contrast, the LoRA + augmentation model requires 625 iterations to achieve a similar equilibrium but with lower overall performance across all metrics. This demonstrates that Meta-LoRA not only accelerates learning but also outperforms standard LoRA in effectiveness.
To assess the impact of augmentation, we compare the performance of LoRA and LoRA + augmentation models. While both models follow a similar trend throughout training, the key distinction lies in their initial R-FaceSim performance at 250 iterations. The higher starting score of the augmented model suggests that augmentation primarily enhances the learning process in the earliest training stages, leading to a more meaningful and stable convergence.
11 Qualitative Results
Figure 6: Qualitative comparisons for an extended list of baselines. See Section 4.5 for explanations.
Figure 6 demonstrates extended comparison for a diverse set of inputs and reference images.
12 Baseline Model Details and Implementation Settings

### v3

@@ 4.2 Experimental setup
Baseline models. We compare our method against three state-of-the-art subject-driven image generation models: InstantID instantid_paper , PhotoMaker photomaker_paper , and PuLID pulid_paper . To ensure a fair comparison, we align evaluation conditions wherever applicable. We also benchmark against a standard LoRA baseline, both with and without augmented training data. Additional details on the baselines, including diffusion backbones and other implementation specifics, are provided in the Supplementary Material.
[…]
@@ 4.3 Evaluation metrics and R-FaceSim
Following recent studies hyperdreambooth_paper ; instantid_paper ; pulid_paper ; instantbooth_paper ; dreambooth_paper , we use two key metrics to evaluate performance: CLIP-T to measure the prompt following ability and the newly proposed Robust Face Similarity (R-FaceSim) to analyze the face similarity. The details of CLIP-T is given in the Supplementary Material.
The commonly used FaceSim metric aims to evaluate how well the generated images maintain the subject’s identity based on the cosine similarity between the reference and generated images. The facial embeddings are obtained using the backbone of a face recognition model. However, we observe two major limitations with it. First, facial recognition methods primarily learn to distinguish individuals, therefore, their embeddings rather tend to lack some of the fine-grained identity details, also observed by facesim_critic ; hyperdreambooth_paper ; kung2024face in various contexts. Second, commonly the same reference images are used for both personalization and face similarity evaluation. However, in feed-forward personalization models, e.g. InstantID instantid_paper and PuLID pulid_paper , we observe a bias towards generating images with the same pose and/or gaze as the reference images. This phenomenon indicates an undesirable limitation of those models. It also tends to artificially inflate the face-similarity scores, because the face recognition embeddings are sensitive to having the same pose or gaze.
To address the former issue, we provide a detailed discussion of the qualitative results. To address the latter one, we propose a novel face-similarity metric that we call Robust Face Similarity (R-FaceSim). R-FaceSim is computed by excluding the reference image (used for fine-tuning or output-conditioning); we instead compare each generated image to other real images of the same person (with different poses) and average the cosine similarity. This yields a face-similarity score not inflated by the exact reference image. We encourage future research to adopt this approach to improve the inflated scores from pose copying as acknowledged in lcm_lookahead_paper .
@@ 4.4 Quantitative Results
Figure 3 provides a quantitative comparison of Meta-LoRA against the previously mentioned baselines on the ’female’ partition of Meta-PHD test set. Our evaluation focuses on two key aspects: CLIP-T, which assess prompt adherence; and R-FaceSim, which measures identity preservation.
Figure 3: Illustration of the metric score trends for Meta-LoRA and the standard LoRA lora_paper models on the Meta-PHD dataset, focusing on the ’female’ class. Left: A plot of R-FaceSim versus CLIP-T scores for all LoRA and Meta-LoRA models trained with ranks $\in$ 1, 2, 4, 8, 16 and across 250, 375, 500, 625, and 750 training iterations. The polynomial fit curves per each model are also shared for better comparison. Right: R-FaceSim and CLIP-T scores plotted against iteration count for rank-1 LoRA and Meta-LoRA models. These results are compared with state-of-the-art models from the literature, including PuLID pulid_paper , InstantID instantid_paper , and PhotoMaker photomaker_paper . To ensure a fair comparison, we use the FLUX.1-dev version of PuLID flux_model , which is consistent with the Meta-LoRA training setup. InstantID and PhotoMaker are based on SD-XL sdxl_paper and SD-XL Lightning sdxl_lightning , as public implementations of these models are not currently available for FLUX.1-dev.
Base models are typically trained for broad, general-purpose tasks, so fine-tuning them for identity personalization can compromise their prompt following ability—often leading to lower CLIP-T scores. Conversely, overfitting to a reference identity can improve R-FaceSim but at the expense of generalization. Striking a balance between these two objectives remains a challenge for existing identity personalization methods. For example, InstantID instantid_paper and PhotoMaker photomaker_paper tend to favor one over the other: PhotoMaker achieves higher CLIP-T scores but compromises identity preservation, while InstantID emphasizes facial similarity at the cost of broader alignment. In contrast, PuLID pulid_paper and our proposed Meta-LoRA demonstrate a more effective balance between identity fidelity and generalization.
On Figure 3 (Left), the training results of Meta-LoRA occupy a higher position on the R-FaceSim vs. CLIP-T plot compared to all other models. Its polynomial trendline outperforms the rest, with LoRA+Augmentation (Aug) and PuLID being the closest in performance. While LoRA+Aug achieves results comparable to PuLID, the default LoRA lags behind all competitors, underscoring the value of data augmentation. Moreover, Meta-LoRA exhibits more stable R-FaceSim values across varying CLIP-T scores, unlike LoRA+Aug. This suggests that Meta-LoRA trains more consistently and learns facial features faster, as evidenced by higher CLIP-T scores in earlier training stages.
On Figure 3 (Right), the effect of iteration count on CLIP-T and R-FaceSim is shown. The plots reveal an inverse relationship between the two metrics. Meta-LoRA captures facial features more quickly than other LoRA variants, highlighting the efficiency of its meta-learned components in reducing adaptation time. However, this comes at the cost of slightly lower text-image alignment performance when compared at the same iteration. Still, between iterations 375 and 500, Meta-LoRA outperforms PuLID on both metrics, suggesting that this range offers a more optimal balance between identity preservation and text alignment. Additional comparisons are given in the Supplementary Material.
@@
Another key difference lies in the scale of pretraining data and the use of additional training losses. Existing models are trained on substantially larger datasets—our dataset constitutes only 0.035% of InstantID’s, 1.4% of PuLID’s, and 18.75% of PhotoMaker’s—and none of these studies share their training data publicly. Furthermore, PuLID incorporates a face-centric identity loss, while InstantID embeds facial features directly into its network. In contrast, Meta-LoRA achieves comparable or superior performance without relying on such explicit facial losses or features, instead employing a simpler and more efficient training strategy.
Figure 4: Qualitative comparisons. Left: Meta-LoRA shows better ID preservation while keeping prompt alignment. Meta-LoRA keeps the pose natural (1st column), promotes identities effectively (2nd and 3rd columns), and follows the given style (4th column). Right: Comparisons across different seeds. For each different seed, Meta-LoRA successfully generates distinct variations, while also preserving the prompt alignment. Details are best viewed on screen, zoomed-in.
@@ 4.5 Qualitative Results
To evaluate visual fidelity and prompt adherence, we qualitatively compare Meta-LoRA against baselines pulid_paper ; photomaker_paper ; lora_paper ; instantid_paper , focusing on female samples due to the larger Stage-1 training dataset (male subject analysis is in Supplementary Material).
Figure 4 (left) underscores Meta-LoRA’s superior ability to preserve nuanced facial characteristics while robustly integrating prompt details. It uniquely achieves a compelling balance between identity retention and scene coherence, producing visually faithful and textually aligned results. In contrast, standard LoRA struggles with consistent identity preservation despite following prompts. InstantID often replicates the reference image’s pose, which, while beneficial for some prompts, significantly curtails its adaptability to diverse viewpoints or styles and can artificially inflate naive similarity metrics. PhotoMaker captures general likeness but often fails to render fine-grained facial details, especially when prompts demand significant contextual or stylistic shifts.
Figure 4 (right) examines robustness to seed variation. Meta-LoRA achieves a desirable equilibrium between seed-dependent output diversity and consistent identity fidelity. Unlike standard LoRA, which can overfit to incidental background features from the reference image, or PuLID, which shows limited variation across seeds and sometimes falters in prompt-driven facial modifications, Meta-LoRA generates diverse yet accurate outputs. This demonstrates its strong capability to explore varied creative outputs while maintaining high fidelity to both the subject’s identity and the textual prompt.
@@
In essence, qualitative assessments strongly corroborate Meta-LoRA’s quantitative advantages. It consistently produces images that are accurate to the subject’s identity and the given prompt, preserving distinctive facial features while offering meaningful creative variation. This positions Meta-LoRA as a highly effective solution for applications demanding both strong identity preservation and versatile prompt adherence.
@@ 5 Conclusion
In this paper, we present Meta-Low-Rank Adaptation (Meta-LoRA), a versatile three-layer LoRA architecture designed to enhance identity personalization in text-to-image generative models. Meta-LoRA employs a two-stage training strategy: the first stage learns generic LoRA Meta-Down layers applicable across multiple identities, while the second stage fine-tunes the LoRA Mid and LoRA Up layers for a specific identity using only a single reference image. By integrating domain-specific priors and optimizing feature learning across diverse identities, Meta-LoRA significantly improves identity preservation, robustness, computational efficiency, and adaptability.
To evaluate our approach, we introduce Meta-PHD, a novel identity personalization dataset used exclusively for testing, and assess Meta-LoRA’s performance using this dataset. Our results demonstrate that Meta-LoRA achieves state-of-the-art (SOTA) performance or performs comparably to existing SOTA methods. Moreover, qualitative evaluations highlight Meta-LoRA’s ability to generate identity-personalized images while effectively maintaining prompt adherence. Future research directions include the incorporation of recent LoRA variants gao24fourierft ; borse2024foura ; yang24vblora ; zhang2023adalora ; kopiczkovera24veralora into the Meta-LoRA framework and the exploration of the proposed method’s applicability to other model customization tasks.
@@ Acknowledgments
Appendix A Implementation Details
Meta-LoRA is implemented using the ai-toolkit111 https://github.com/ostris/ai-toolkit framework. Training and evaluation are performed on a single NVIDIA RTX A6000 GPU with 48GB of VRAM (see Section C for details). Under this setup, Stage-1 training takes approximately 7 days (13 seconds per iteration for a total of 50,000 iterations), while Stage-2 training takes around 18 minutes (2.8 seconds per iteration for 375 iterations).
For all training runs, we use a learning rate of $0.0004$, set the total number of Stage-1 iterations ($q_{total}$) to 50,000, bucket size ($q_{bucket}$) to 2,500, warm-up iterations ($q_{warm\text{-}up}$) to 1,000, and Stage-2 iterations ($q_{st2}$) to 375. To analyze performance across configurations, we evaluate both Meta-LoRA and the default LoRA at different ranks $\in\{1,2,4,8,16\}$ and checkpoints $\in\{250,375,500,625,750\}$. Batch sizes are set to 4 for Stage-1 and 1 for Stage-2. Training is performed using the standard L2 loss, and we use the AdamW optimizer [19].
Appendix B Meta-Training and Personalization Algorithms
@@
Appendix C Baseline Model Details and Implementation Settings